A cylinder which is in a horizontal position contains an unknown noble gas at 4.84 x 104 Pa and is sealed with a massless piston. The piston is slowly, isobarically moved inward 0.189 m, while 1.95 × 104 J of heat is removed from the gas. If the piston has a radius of 0.338 m, calculate the change in internal energy of the system. Number
Answer #1

Solution,

dU = dQ - dW

dW = P dV = (48400) x pi x (0.338)^2 x 0.189 = 3283.15 J

dU = - 19500 - 3283.15 = - 22783.15 J

Comment in case any doubt please rate my answer .....